OpenStudy (anonymous):

Solve the given differential equation by separation of variables x dy/dx = 4y Please show work

OpenStudy (unklerhaukus):

\[x\frac{\mathrm dy}{\mathrm dx} = 4y\] separate the variables \[\frac{\mathrm dy}y=4\frac{\mathrm dx}x \] now integrate

OpenStudy (unklerhaukus):

\[\int\frac{\mathrm dy}y=4\int\frac{\mathrm dx}x\]

OpenStudy (anonymous):

Thank you to the both of you!

OpenStudy (anonymous):

lol nvm one person

OpenStudy (unklerhaukus):

ii can be two persons

OpenStudy (unklerhaukus):

NB: dont for get the constant of integration